An item costs $900, has a scrap value of $50, and a useful life of five years. The linear equation relating book value and number of years is:

Answer by stanbon(48568)

An item costs $900, has a scrap value of $50, and a useful life of five years. The linear equation relating book value and number of years is:
-----------------------
You have two points: (0,900) and (5,50)
slope = (50-900)/(5-0) = -850/5 = -170
-------------------
intercept = 900
------------------
EQUATION:
book value = -170x + 900 where x is the number of years from 0 to 5
----------------------
Cheers,
Stan H.